OpenStudy (kaylak12345):

geo help a diameter has endpoints -10, -2 and 4,6 what is the radius center and equation of the circle heres an example question help me through it

OpenStudy (kaylak12345):

please help me this is too hard

OpenStudy (kaylak12345):

hurry people i have things to do and im stuck here

OpenStudy (welshfella):

first find the midpoint of the line joining the 2 given points can you do that?

OpenStudy (kaylak12345):

is it -3,2

OpenStudy (welshfella):

yes Thats the coordinates of the center Now use the distance formula to fin diameter of the circle then you'll get the radius

OpenStudy (welshfella):

diameter = sqrt (6 - -2)^2 + ( 4 --10)^2

OpenStudy (kaylak12345):

64+196 square root

OpenStudy (kaylak12345):

about 16.1

OpenStudy (welshfella):

16.12 so radius = 8.06 now you plug in these values into general equation for a circle//;- (x - a)^2 + (y - b)^2 = r^2 where center = (a,b) and r = radius

OpenStudy (welshfella):

a = -3 , b = 2 and r^2 = 64.96

OpenStudy (kaylak12345):

(x+3)^2+(y-2)^2=64.96

OpenStudy (welshfella):

right!

OpenStudy (kaylak12345):

yay thank you

OpenStudy (welshfella):

yw

OpenStudy (sshayer):

you can also write the eq. of circle by the extremities of diameter. \[\left( x-x _{1} \right)\left( x-x _{2} \right)+\left( y-y _{1} \right)\left( y-y _{2} \right)=0 \] general eq. of circle is \[x^2+y^2+2gx+2fy+c=0\] where center is (-g,-f) and \[radius=\sqrt{g^2+f^2-c}\]